OpenStudy (anonymous):

the math club has 57 members who are sixth and seventh graders.There twice as many sixth graders as seventh graders. How many sixth graders are there?

OpenStudy (anonymous):

I did 57 divided by 3. This gave me 19. 19 + 19 is 38. So I came up with there are 38 6th graders and 19 seventh graders. Hope I helped!

OpenStudy (anonymous):

thanks....so how do I teach this to a sixth grader, so that he'll be able to do the next problem? Home schooling.

OpenStudy (anonymous):

Ahh the joys of homeschooling. I have no clue but I'll try to think of something that makes sense. All kids learn differently but how I learned way That you had 3 groups but you merged two. So say it's like a video game! You have 3 teams but only enough flags for 2 teams so you need to combine two of those teams to make one. Does this make some sense? I hope so.

OpenStudy (anonymous):

lol......I like it!! thanks

OpenStudy (anonymous):

np!! Good luck with the homeschooling!!

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

:D If you need any more help feel free to tag me and I'll try to help :) @ohlookidontunderstandanotherthing

OpenStudy (anonymous):

I'll take you up on that. thanks

OpenStudy (anonymous):

hello friend,I have another math qusetion for you.The population of Medfield is 7000. that is 8/25 of the population of capital city. What is Capital City's population? I got 2240. I did 7000 x 8/25=56000/25=2240. Am I wrong? I have to be because that would make Medfield bigger than Capital City. smh

OpenStudy (anonymous):

Hello friend! Apologies it's taken me so long to answer, I don't have open study set to alert me :P Heres what I did 7000=8/20 of the population so I did 7000 divided by 8=825. so 1/20=825 people. Still with me? I hope so. Now I did 825 X 20 = 16500 I got The population of Capital City is 16500. Sorry again it took so long to answer!!

OpenStudy (anonymous):

No apologies necessary! ok. Why did you do 7000=8/20 instead of 7000=8/25 ? then 7000 x 25= 175000. then 17500 divided by 8 =21875....21875 is what I got. I'm not sure if that's right.

OpenStudy (anonymous):

I did it wrong! My brain was seeing 20 instead of 25. Lemme try this again 825X25=20625 But remember that this is only 8/25! Meaning that this is only 25/100 of the answer. Add this four times and get 82500. I believe the population is 82,500 people. Sorry I was rushing a bit earlier!!